OpenStudy (anonymous):

A family has a total income of $32,400. Of this 1/4 spent for food, 1/5 for housing, 1/10 for clothing, 1/9 for savings, and 1/4 for taxes and the rest for other expenses. How much is spent for housing?

OpenStudy (anonymous):

They say that \(\frac{1}{5}\) is used for housing so you are taking: \[32400\times\frac{1}{5}=6480\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No prob :)